Skip to Main Content
PTAGIS logo link back to home page

Interrogation Site Metadata

Search by typing either a site name, site code, organization, or site type.

ESS - EFSF Salmon River at Parks Cr.

Status:   Active - Operations Unspecified
Site Type: Instream Remote Detection System  |  Organization: Biomark
General Operational Period: Year Round
Start Date: 3/26/2009 12:00 PM

This is an in-stream interrogation system consisting of two separate antenna arrays, 790 meters apart, on the East Fk South Fk Salmon River (rkm 21) near Parks Creek. The upstream and downstream arrays consist of four flat plate antennas each, spanning 24 meters across the low-water channel.

Current Operations

Last File Loaded: ESS-2021-295-P-00H.xml
Date Last File Loaded: 10/22/2021 8:50 PM
Date Last File Opened: 10/22/2021 5:16 PM
Date Last File Closed: 10/22/2021 7:04 PM
Last Observation Tag: 3DD.003D6EDF60
Last Observation Date: 10/22/2021 6:51 PM
Detection Summary: Last Day's Observations


Primary Contact:

Kyle Meier

Data Steward:

Brett Turley


Coordinates: 44.956205, -115.533150  |  River Km: 522.303.215.060.021

Operational History

Operational history describes the period the site was interrogating for PIT tags. The comments provide anecdotal information about the operation of the site during this period. Site operational periods are updated typically at the end of the calendar year.

Year Start Date End Date Comments
2020 1/1/2020 12:00 AM 11/24/2020 12:00 AM Start and end dates based on timer tags. See event logs for additional operational information.
2019 1/1/2019 12:00 AM 12/31/2019 12:00 AM Start and end dates based on timer tags.
2018 1/1/2018 12:00 AM 1/1/2019 12:00 AM Start and end dates based on timer tags.
2017 1/1/2017 12:00 AM 1/1/2018 12:00 AM Start and end dates based on timer tags.
2016 1/1/2016 12:00 AM 1/1/2017 12:00 AM Start and end dates based on timer tags.
2015 1/1/2015 12:00 AM 1/1/2016 12:00 AM
2014 1/1/2014 12:00 AM 1/1/2015 12:00 AM
2013 1/1/2013 12:00 AM 1/1/2014 12:00 AM
2012 1/1/2012 12:00 AM 1/1/2013 12:00 AM
2011 1/1/2011 12:00 AM 1/1/2012 12:00 AM The upper array (A0) lost power for 25 hours starting on March 1. Data from April 15-22 was lost due to data logger malfunction.
2010 1/1/2010 12:00 AM 1/1/2011 12:00 AM
2009 3/26/2009 12:00 PM 1/1/2010 12:00 AM First year of operation. The ESS A0 array (upstream) was installed during the week of March 23, 2009. The array consists of 4-20 ft Biomark pass-by antennas. Monitoring began at approximately noon on March 26, 2009. The ESS B0 array (downstream) became operational on August 6, 2009.


Configuration metadata includes details about the equipment used at the site to detect and record PIT tags. A new configuration sequence record is created when equipment is added or removed from the site and the period of time it was active. A configuration sequence without an End Date is currently active.

Select Configuration Number:
Start Date:  3/1/2009
End Date:  ----

Equipment History

Select a type to view the history of equipment installed at this site. A missing End Date implies the information about the equipment is current.

Select Equipment Type:
Equipment Code Description Start Date End Date

No Photos

Event Log Summary

Event Date Event Type Comments
9/17/2020 2:06:00 PM General East Fork South Fork Salmon (ESS B0) antenna B2 was showing extremely high noise. Antenna cable was replaced on antenna B2 resolving the high noise. ESS A0 and ESS B0 PIT arrays are running continuously without disruption through September 21, 2020.
11/25/2019 10:11:34 AM General Sites ESS A0 and ESS B0 have been operating continuously without disruption through November 25, 2019.
11/30/2017 7:44:28 AM General ESSA0 array antenna A2 cable was replaced on September 13, 2017 due to high and variable noise. The cable replacement did not solve the problem. ESSB0 array antenna cable B2 came unplugged during high spring flows. The cable was reattached and secured on September 13, 2017. ESSA0 and ESSB0 has run continuously without disruption through November 30, 2017.
9/19/2016 9:08:00 AM General ESS B0 antenna B2 was replaced with a new antenna on 9/19/2016 due to low current. ESS A0 and ESS B0 have run without disruption through 11/16/16.
10/23/2013 9:31:32 AM General The site ESS A0 ran out propane fuel on 9/20/13 resulting in an interruption in data collection. The propane tank was filled and normal operation resumed on 9/25/13. The site has been running continuously without disruption through 10/23/13. The site ESS B0 has been running continuously without disruption through 10/23/13

Archived Event Logs